Accelerator Update for April 15, 2009 - April 17, 2009
Wednesday April 15
The day shift began with Operations monitoring the Tevatron (TeV) store 6978 (8.24E31) and the Recycler stash (292.1E10), with the Antiproton Source stacking (26.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
Operations established store 6979 by 1:25 PM with an average initial luminosity of 32E31.
The evening shift maintained HEP support operations.
Thursday April 16
The midnight shift began with Operations monitoring store 6979 (10.1E31) and the stash (270.3E10), with Pbar stacking (32.4E10), and with NuMI and MiniBooNE taking beam.
Operations terminated store 6979 at 4:30 AM and HEP shot setup followed.
Operations established store 6980 around 5:55 AM with an initial luminosity of 34.1E31.
Booster and MI personnel tuned their machines for most of the day shift.
A MI RF station (MIRF5) tripped off. Technicians replaced a modulator resistor during the afternoon.
Operations disabled the MTest beam at 6 PM.
Operations terminated store 6980 at 11 PM and put the TeV into shot setup.
Friday April 17
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (11.9E10), with Pbar stacking (44.5E10), and with NuMI and MiniBooNE taking beam.
Operations established store 6981 at 12:48 AM with an initial luminosity of 31.9E31.
The Neutron Therapy Facility reported at 7:25 AM that they were tuning up in preparation to treat one patient later in the morning.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 04/17/2009 email Fermilab|
|Security, Privacy, Legal|